2.42 Show that the circuit in Fig. P2.42 is all pass. Specifically: (a) Show that its gain is unity at all frequencies.(b) Find the frequency wo, in terms of R and C, for which if vi(t) = cos(wot) then v2(t) = sin(wot). This is called a 90-degr

Do without using Laplace
Answer #1

(4) -loswot Va I + Valt) - TV Write in phosor wilt) = Coswt) V = 1 LO & V₂ H) = Volt) - Vy(t) write into phasor form V = Va Vhet V-V-7 - Pyjoe = Vo - Vo = 14 l RC - WRC Ve=V(jakso Hlin) 12 = 1-jWRC 1 1+jwc imput Gain 1 + (WRC) ² 1+(WRC)2 V₂ = outputVC4) = Hiw) | www.Smw.t) =/(ww.cos(wat - 909) where $ = - 90=/ Hiw|ww. compare equation on she H(jw) / w-5-90 = tan 1 - WRC)